3. Rambo Companys owners equity equals % of the companys total assets. The companys liabilities are $120,000. What is the
Answer #1

Assets = Liabilities + Owner's Equity

Let, Assets = 'X' therefore, Owner's Equity = 1/2 * X Liability = 120000

Asset = Liability + Equity X = 120000 + (1/2 * X)     X - (X/2) = 120000     X/2 = 120000     X = 2*120000    X = $240000

Asset = $240000

Equity = 1/2 * 240000

Equity =$ 120000